OpenStudy (anonymous):

Without a calculator how do i solve log16^8=(x+1) move to exponent form..

OpenStudy (whpalmer4):

what is the base of the logarithm here?

OpenStudy (anonymous):

16 @whpalmer4

OpenStudy (whpalmer4):

so you have \[\log_{16}8 = x+1\]?

OpenStudy (anonymous):

yes @whpalmer4

OpenStudy (whpalmer4):

remember, if \[\log_b u = a\]then \[b^a = u\]

OpenStudy (anonymous):

so it would be 16(x+1)=8

OpenStudy (anonymous):

? @whpalmer4

OpenStudy (whpalmer4):

Uh, put a "^" in there if you want to indicate an exponent

OpenStudy (anonymous):

16^(x+1)=8

OpenStudy (whpalmer4):

Do you mean \[16(x+1) = 8\]or\[16^{x+1} = 8\]

OpenStudy (whpalmer4):

Right. So how can you solve that?

OpenStudy (anonymous):

im not sure lol

OpenStudy (whpalmer4):

16 and 8 are both powers of 2, aren't they?

OpenStudy (anonymous):

yeah

OpenStudy (whpalmer4):

what are they, expressed as powers of 2?

OpenStudy (whpalmer4):

\(2^8 = 16\)?

OpenStudy (whpalmer4):

2*2=4 2*2*2=8 2*2*2*2=16

OpenStudy (anonymous):

okay

OpenStudy (whpalmer4):

so \[16=2^4\] and \[8=2^3\]right?

OpenStudy (anonymous):

yup

OpenStudy (whpalmer4):

Okay, do you remember that \[(ab)^n = a^nb^n\]

OpenStudy (anonymous):

yeah

OpenStudy (whpalmer4):

Okay, so we can rewrite our equation as \[(2*2*2*2)^{x+1} = 2^3\]

OpenStudy (anonymous):

got it

OpenStudy (whpalmer4):

what's your final answer?

OpenStudy (anonymous):

oh was that the final step before the answer?

OpenStudy (whpalmer4):

yeah, you still want to find \(x\), right?

OpenStudy (anonymous):

yeah but if i were to factor that again wouldnt it just give me the same equation as before "16^x+1=8"

OpenStudy (whpalmer4):

no, use that property I just showed you:\[(ab)^n = a^nb^n\]

OpenStudy (anonymous):

whats the little symbols next to this (ab)? say its tiny

OpenStudy (whpalmer4):

oh, try this: \[\huge (ab)^n = a^nb^n\]

OpenStudy (whpalmer4):

Here, I'll just write it out for you: \[\huge (2*2*2*2)^{x+1} = 2^3\]\[\huge 2^{x+1}2^{x+1}2^{x+1}2^{x+1} = 2^3\]

OpenStudy (whpalmer4):

And remember that \[\huge x^nx^m = x^{n+m}\]

OpenStudy (whpalmer4):

So that gives us \[\huge 2^{4(x+1)} = 2^{3}\]\[x=\]

OpenStudy (anonymous):

i got -1/4 @whpalmer4

OpenStudy (whpalmer4):

Very good!

OpenStudy (anonymous):

Thank you so much, ill give you a medal :)

OpenStudy (whpalmer4):

1000 medals and $3.50 gets you a cup of coffee at Starbucks :-)

OpenStudy (anonymous):

interesting, how many medals do you have?

OpenStudy (whpalmer4):

(they are always appreciated, even if worthless!)

OpenStudy (whpalmer4):

2093, apparently. If you click on someone's name, the little window that pops up gives you a link to click to see their entire profile, which shows how many fans, how many medals, questions answered/asked, shoe size, etc.

OpenStudy (whpalmer4):

You could also have solved the last bit of this problem by noticing that \[(2^4)^{x+1} = 2^3\]Remember that \[(u^a)^b = u^{a*b}\]so we have \[2^{4*(x+1)} = 2^3\]

OpenStudy (anonymous):

wow thats so awesome, and you just received another fan hehe thanks :)